OpenStudy (anonymous):

Using the equation y = 3x + 7, for every two units the graph moves left, how many units does the graph move up?

OpenStudy (anonymous):

if you move the graph to the left then the equation for it is y=3(x+1)+7 = 3x + 3 + 7 = 3x+10.. Because 10 is 3 more than 7, then the graph must move 3 units up.

OpenStudy (anonymous):

oh that's wrong, my example deals with one unit to the left. If you want two units to the left then try this: y=3(x+2)+7 = 3x + 13. 13-7 = 6 so the graph moves 6 units up for every two units it moves to the left.